OpenStudy (anonymous):

A train on a mountain railway is carrying 200 people of average mass 70 kg up a slope at an angle of 30° to the horizontal and at a speed of 6.0 m s–1. The train itself has a mass of 80 000 kg. The percentage of the power from the engine which is used to raise the passengers and the train is 40 %. What is the power of the engine? A 1.1 MW B 2.8 MW C 6.9 MW D 14 MW

OpenStudy (somy):

what is the answer supposed to be? im getting D

OpenStudy (somy):

ouh no i got it its C

OpenStudy (anonymous):

It says C.

OpenStudy (anonymous):

Oh, I got it too. 6.9 MW. I don't know why I was making it so complicated. Just remembered that I can use the P=Fv formula here. >_<

OpenStudy (somy):

yeah thats what i did

OpenStudy (somy):

its always like that some we tend to make it more complicated then it already is lol

OpenStudy (anonymous):

Yea. :/ Okay guys help me with this one.

OpenStudy (anonymous):

great I did that completely on my own

OpenStudy (anonymous):

no I mean first question lol

OpenStudy (anonymous):

Good. xD

OpenStudy (somy):

okaaay lets seee

OpenStudy (anonymous):

I'm getting C.

OpenStudy (anonymous):

Oh, I just got what I was doing wrong. I was simply multiplying the masses with distances. HOW STUPID of me! -_-

OpenStudy (anonymous):

If you multiply C i.e 0.29 with 9.81, you get D 2.8 Nm. What is wrong with me today? :'( Earlier, I was using F=mgh.

OpenStudy (somy):

how did u do it???

OpenStudy (anonymous):

Sum of clockwise moments= Sum of anticlockwise moments

OpenStudy (somy):

that's what i did but i got a weird answer, can u take a pic of ur working plz?

OpenStudy (anonymous):

\[6g \times 0.15 + X = (3g \times 0.35) + (1.4g \times 0.1)\]

OpenStudy (anonymous):

X is the torque we have to find.

OpenStudy (anonymous):

guys why do we take the component of the speed not the component of the force ?

OpenStudy (anonymous):

I don't think you need to worry about that since you're using the formula P= Fv then that means you take the value of the speed from the triangle you draw and then multiply that v with the F

OpenStudy (anonymous):

ohright thank you so much do you guys take AS level as well ?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

and no prob :)

OpenStudy (anonymous):

Wow , its a pleasure .. :)

OpenStudy (anonymous):

WAIT , but why do we divide by 0.4 shouldnt it be time 0.4 ?

OpenStudy (somy):

@Lama97 where did u bring 0.4 from?

OpenStudy (anonymous):

Im talkig about the first question... like do we multiply by 40\100 or divide >

OpenStudy (somy):

i'd suggest you to draw what you are reading u'll get a triangle it's easier to understand what is going on